OpenStudy (anonymous):

i have a problem. perform the operation, leave results in factored form. x/x^2-7x+6 - x/x^2-2x-24 i have the denominator right but i keep getting 10x instead of 5x for the numerator, help

OpenStudy (anonymous):

\[\frac{x}{(x - 6)(x -1)} - \frac{x}{(x-6)(x +2)}\]

OpenStudy (anonymous):

\[\frac{x^2 +2x -x^2 +x}{(x -1)(x - 6)( x +2)}\]

OpenStudy (anonymous):

now you can have the ans

OpenStudy (anonymous):

the second fraction, seems strange. wouldn't the product of x-6 and x+2 give you x^2-4x -12

OpenStudy (anonymous):

I found my mistake, thanks for the help